OpenStudy (anonymous):

do 6 : 15 and 30 : 80 make a proportion? if that's how I'd state it...

OpenStudy (anonymous):

Simplify them Imagine that they are a fraction, like this: \[\frac{ 6 }{ 15 }=\frac{ 2*3 }{ 3*5 }=\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

Do the same with 30 : 80 and if you get the same result than they are in proportion.

OpenStudy (anonymous):

Thank you for your help!